OpenStudy (anonymous):

what are the 2 solutions to 5x squared +8x -4=0

OpenStudy (anonymous):

0.4 or -2

OpenStudy (anonymous):

You can give medals by clicking 'Best Response'.

OpenStudy (triciaal):

5x^2 + 8x -4 = 0 what are the factors of 20 that add up to 8 use 10,-2 5x^2 +10x -2x-4 = 0 5x( x+ 2) -2(x+2) = 0 (5x-2)(x + 2) = 0 5x-2 = 0; x + 2 =0 answers were given above

OpenStudy (anonymous):

the answer said -2\5. or -2 is that right

OpenStudy (triciaal):

i got 2/5 positive double check

OpenStudy (anonymous):

o u rite

OpenStudy (triciaal):

ok do you understand the steps?

OpenStudy (anonymous):

yes i jus needed some input

OpenStudy (triciaal):

ok

OpenStudy (anonymous):

what is the square root of 289

OpenStudy (triciaal):

17